# Equal Contribution

### v4

@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved AI capabilities OpenAI (2023). LLMs have access to a vast amount of training data, allowing them to acquire a substantial amount of knowledge, ranging from common-sense to specific domains, such as math and science. Moreover, recent papers have argued that complex behaviors, such as writing code, generating long stories, and even some reasoning capabilities, can emerge from such large-scale training Wei et al. (2022); Rozière et al. (2023); Zhao et al. (2023); Yao et al. (2023). Impressively, LLMs can use their knowledge to generate plausible explanations given observations and even reason on counterfactuals (Wang et al., 2023; Li et al., 2023; Kıcıman et al., 2023). Given these emergent abilities to understand and learn relationships from textual corpora, a natural question arises: Can LLMs reason about causal relationships?
@@
Traditional causal discovery frameworks Bello et al. (2023); Zheng et al. (2018); Meek (2023) rely on numerical observations and use statistical or continuous optimization based methods to unravel confounding and causal factors. While such rigorous numerical and abstract reasoning abilities remain challenges for LLMs (Gendron et al., 2023; Xu et al., 2023; Imani et al., 2023), recent works have shown that LLMs are capable of leveraging their internal knowledge to perform commonsense and simple logical as well as causal reasoning Wei et al. (2022); Zhao et al. (2023); Yao et al. (2023); Kıcıman et al. (2023). Inspired by these results, the proposed framework utilizes LLMs to respond to causal queries solely based on metadata, such as variable names and descriptions, without accessing the numerical observations. This is similar to the metadata and knowledge-based reasoning that is done by human domain experts when they construct causal graphs.
Previous works Kıcıman et al. (2023); Choi et al. (2022); Long et al. (2023b) have shown that LLMs are capable of inferring pairwise causal relationships given only metadata. However, extending this pairwise framework to full graph discovery presents additional challenges. One straightforward approach which previous works use is to perform pairwise queries over all possible pairs of variables. However, this method is not scalable since it requires ($n\choose 2$ total queries where $n$ is the number of variables. Thus, the number of queries required grows quadratically with the number of causal variables. As a result, the experiments in these works are limited to very small causal graphs.
@@ 2 Related Work
Causal graph discovery has been extensively studied in many settings Zanga & Stella (2023). Until recently, most proposed methods were statistical methods that used observational data to infer causal relationships between variables. Some examples of these are Greedy Equivalence Search (GES) Meek (2023), the Peter-Clark Algorithm (PC) Spirtes & Glymour (1991), Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) Zheng et al. (2018), and Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) Bello et al. (2023). Recently, many works have proposed to use LLMs for causal graph discovery Kıcıman et al. (2023); Choi et al. (2022); Long et al. (2023b). These LLM-based methods do not rely on observational data. Instead, they utilize metadata associated with the input variables, specifically variable descriptions in natural language. This approach closely resembles how human experts construct causal graphs using their domain knowledge.
@@ C. There is no causal relationship between <A> and <B>
To extend to full graph discovery, Kıcıman et al. (2023) propose to perform pairwise queries over all possible pairs. However, this method is not scalable since it requires ($n\choose 2$ $\in O(n^{2})$ total queries, where $n$ is the number of variables, so they are only able to perform experiments on a small causal graph with 12 variables.
@@ 3.2 Adding Statistics to the Prompt
$G\leftarrow\{\}$triangleright Create an empty graph to store the result.
$frontier,visited\leftarrow I(p_{\theta},X)$triangleright With initialization prompt.
while $frontier$ is not empty do
@@
     for $node$ in $E(p_{\theta},G)$ do $\triangleright$ Expand with expansion prompt.
@@
         if $node$ not in $frontier\cup visited$ then
@@ 4.1 Experimental Setup
    Asia Lauritzen & Spiegelhalter (1988). The Asia causal graph is a small causal graph that models lung cancer in patients having potentially recently visited Asia. The ground-truth graph has 8 nodes and 8 edges. Some examples of variables in the graph are Visit to Asia, Smoking, Bronchitis, and Lung Cancer.
@@
We compare the proposed method, as well as the proposed method plus observational statistics from 1000 and 10000 samples, against the numerical methods of Greedy Equivalence Search (GES) Meek (2023), the Peter-Clark Algorithm (PC) Spirtes & Glymour (1991), Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) Zheng et al. (2018), and Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) Bello et al. (2023), as well as against pairwise queries, and pairwise queries plus observational statistics from 1000 and 10000 samples. We use the implementations of GES and PC from the causal-learn package Zheng et al. (2023), the implementation of NOTEARS from Zheng et al. (2018), and the implementation of DAGMA from Bello et al. (2023). For the numerical methods, we conduct experiments in the low (100 samples), medium (1000 samples), and high (10000 samples) data regimes.
@@ 4.2 Metrics
Following Kıcıman et al. (2023), we also report the Normalized Hamming Distance (NHD). In a graph with $m$ nodes, the NHD is given by $\sum_{i,j=1}^{m}\frac{1}{m^{2}}\mathbbm{1}_{G_{ij}\neq G^{\prime}_{ij}}$, i.e. the number of edges that are present in one graph and not the other, divided by the total number of all possible edges. One problem with the NHD is that since most graphs are relatively sparse, trivially outputting 0 predicted edges achieves a relatively good NHD. For example, for the Child graph, which has 20 nodes and 25 edges, if we output 0 predicted edges then we achieve a NHD of $\frac{1}{(20^{2})}25=0.0625$, which is better than the NHD achieved by all the methods. Intuitively, predicting an edge is much riskier when using this metric, so algorithms that predict more edges are punished. To remedy this and take into account the number of predicted edges, again following Kıcıman et al. (2023), we compute the ratio between the NHD and the baseline NHD of an algorithm that outputs the same number of edges but with all of them being incorrect. The lower the ratio, the better the method does compared to the worst baseline that outputs the same number of edges. We report the NHD, number of predicted edges (NPE), baseline NHD, and ratio between NHD and baseline NHD for all algorithms.
@@ 4.3 Results
On the Neuropathic Pain causal graph, running pairwise queries would require ($221\choose 2$ $=24310$ queries, which is impractical. Similarly, GES requires considering all 24310 possible edges at each step of its search, which is intractable. We thus omit the results for both GES and pairwise queries on the Neuropathic Pain graph. In addition, PC runs out of memory when run on the Neuropathic Pain graph with 10000 samples, so we also omit those results from the table.
@@ 7 Future Work
Another avenue to explore is using advanced prompting strategies like Tree of Thoughts Yao et al. (2023), which are known to significantly improve performance in complex multi-step reasoning tasks like causal graph discovery.
@@ 8 Reproducibility Statement
We provide code to reproduce all experiments in this anonymized repository

### v5

@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly improved AI capabilities OpenAI (2023). LLMs have access to a vast amount of training data, allowing them to acquire a substantial amount of knowledge, ranging from common-sense to specific domains, such as math and science. Moreover, recent papers have argued that complex behaviors, such as writing code, generating long stories, and even some reasoning capabilities, can emerge from such large-scale training Wei et al. (2022); rozière2023code; Zhao et al. (2023); Yao et al. (2023). Impressively, LLMs can use their knowledge to generate plausible explanations given observations and even reason on counterfactuals (Wang et al., 2023; Li et al., 2023; kıcıman2023causal). Given these emergent abilities to understand and learn relationships from textual corpora, a natural question arises: Can LLMs reason about causal relationships?
@@
Traditional causal discovery frameworks Bello et al. (2023); Zheng et al. (2018); Meek (2023) rely on numerical observations and use statistical or continuous optimization based methods to unravel confounding and causal factors. While such rigorous numerical and abstract reasoning abilities remain challenges for LLMs (Gendron et al., 2023; Xu et al., 2023; Imani et al., 2023), recent works have shown that LLMs are capable of leveraging their internal knowledge to perform commonsense and simple logical as well as causal reasoning Wei et al. (2022); Zhao et al. (2023); Yao et al. (2023); kıcıman2023causal. Inspired by these results, the proposed framework utilizes LLMs to respond to causal queries solely based on metadata, such as variable names and descriptions, without accessing the numerical observations. This is similar to the metadata and knowledge-based reasoning that is done by human domain experts when they construct causal graphs.
Previous works kıcıman2023causal; Choi et al. (2022); Long et al. (2023b) have shown that LLMs are capable of inferring pairwise causal relationships given only metadata. However, extending this pairwise framework to full graph discovery presents additional challenges. One straightforward approach which previous works use is to perform pairwise queries over all possible pairs of variables. However, this method is not scalable since it requires $n\choose 2$ total queries where $n$ is the number of variables. Thus, the number of queries required grows quadratically with the number of causal variables. As a result, the experiments in these works are limited to very small causal graphs.
@@ 2 Related Work
Causal graph discovery has been extensively studied in many settings Zanga and Stella (2023). Until recently, most proposed methods were statistical methods that used observational data to infer causal relationships between variables. Some examples of these are Greedy Equivalence Search (GES) Meek (2023), the Peter-Clark Algorithm (PC) Spirtes and Glymour (1991), Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) Zheng et al. (2018), and Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) Bello et al. (2023). Recently, many works have proposed to use LLMs for causal graph discovery kıcıman2023causal; Choi et al. (2022); Long et al. (2023b). These LLM-based methods do not rely on observational data. Instead, they utilize metadata associated with the input variables, specifically variable descriptions in natural language. This approach closely resembles how human experts construct causal graphs using their domain knowledge.
@@ C. There is no causal relationship between <A> and <B>
To extend to full graph discovery, kıcıman2023causal propose to perform pairwise queries over all possible pairs. However, this method is not scalable since it requires $n\choose 2$ $\in O(n^{2})$ total queries, where $n$ is the number of variables, so they are only able to perform experiments on a small causal graph with 12 variables.
@@ 3.2 Adding Statistics to the Prompt
$G\leftarrow\{\}$ $\triangleright$ Create an empty graph to store the result.
$frontier,visited\leftarrow I(p_{\theta},X)$ $\triangleright$ With initialization prompt.
while frontierfrontier is not empty do
@@
  for nodenode in $E(p_{\theta},G)$ do $\triangleright$ Expand with expansion prompt.
@@
   if nodenode not in $frontier\cup visited$ then
@@ 4.1 Experimental Setup
    Asia Lauritzen and Spiegelhalter (1988). The Asia causal graph is a small causal graph that models lung cancer in patients having potentially recently visited Asia. The ground-truth graph has 8 nodes and 8 edges. Some examples of variables in the graph are Visit to Asia, Smoking, Bronchitis, and Lung Cancer.
@@
We compare the proposed method, as well as the proposed method plus observational statistics from 1000 and 10000 samples, against the numerical methods of Greedy Equivalence Search (GES) Meek (2023), the Peter-Clark Algorithm (PC) Spirtes and Glymour (1991), Non-combinatorial Optimization via Trace Exponential and Augmented lagrangian for Structure learning (NOTEARS) Zheng et al. (2018), and Directed Acyclic Graphs via M-matrices for Acyclicity (DAGMA) Bello et al. (2023), as well as against pairwise queries, and pairwise queries plus observational statistics from 1000 and 10000 samples. We use the implementations of GES and PC from the causal-learn package Zheng et al. (2023), the implementation of NOTEARS from Zheng et al. (2018), and the implementation of DAGMA from Bello et al. (2023). For the numerical methods, we conduct experiments in the low (100 samples), medium (1000 samples), and high (10000 samples) data regimes.
@@ 4.2 Metrics
Following kıcıman2023causal, we also report the Normalized Hamming Distance (NHD). In a graph with $m$ nodes, the NHD is given by $\sum_{i,j=1}^{m}\frac{1}{m^{2}}\mathbbm{1}_{G_{ij}\neq G^{\prime}_{ij}}$, i.e. the number of edges that are present in one graph and not the other, divided by the total number of all possible edges. One problem with the NHD is that since most graphs are relatively sparse, trivially outputting 0 predicted edges achieves a relatively good NHD. For example, for the Child graph, which has 20 nodes and 25 edges, if we output 0 predicted edges then we achieve a NHD of $\frac{1}{(20^{2})}25=0.0625$, which is better than the NHD achieved by all the methods. Intuitively, predicting an edge is much riskier when using this metric, so algorithms that predict more edges are punished. To remedy this and take into account the number of predicted edges, again following kıcıman2023causal, we compute the ratio between the NHD and the baseline NHD of an algorithm that outputs the same number of edges but with all of them being incorrect. The lower the ratio, the better the method does compared to the worst baseline that outputs the same number of edges. We report the NHD, number of predicted edges (NPE), baseline NHD, and ratio between NHD and baseline NHD for all algorithms.
@@ 4.3 Results
On the Neuropathic Pain causal graph, running pairwise queries would require $221\choose 2$ =24310=24310 queries, which is impractical. Similarly, GES requires considering all 24310 possible edges at each step of its search, which is intractable. We thus omit the results for both GES and pairwise queries on the Neuropathic Pain graph. In addition, PC runs out of memory when run on the Neuropathic Pain graph with 10000 samples, so we also omit those results from the table.
@@ 7 Future Work
Another avenue to explore is using advanced prompting strategies like Tree of Thoughts (Yao et al., 2023), which are known to significantly improve performance in complex multi-step reasoning tasks like causal graph discovery.
@@ 8 Reproducibility Statement
We provide code to reproduce all experiments in this repository